when thirty times a number is increases by 32, the result is equal to twice the square of the number. Find tne number.
Answer by lwsshak3(11628)

when thirty times a number is increases by 32, the result is equal to twice the square of the number. Find tne number.
***
let x=number
30x+32=2x^2
2x^2-30x-32=0
x^2-15x-16=0
(x-16)(x+1)=0
x=16
number=16